On a mission to a newly discovered planet, an astronaut finds chlorine abundances of 13.85 for and 86.15 for . What is the atomic mass of chlorine for this location?
The mass of is 34.97 . The mass of is 36.97 .
Express your answer in atomic mass units using four significant figures.
1
Expert's answer
2012-10-01T10:05:06-0400
The atomic mass of chlorine for this location is
M(Cl) = M(Cl-35)*x(Cl-35) +
M(Cl-37)*x(Cl-37) = 34.97*0.1385 + 36.97*0.8615 = 36.69